# Effective Polaron Dynamics of an Impurity Particle Interacting with a Fermi Gas

### v1

@@ Appendix A Estimates within the bosonization framework
$=2\kappa k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)^{\ast}$
One makes the identification $\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\equiv c_{\alpha}^{g}(k)$ representing a weighted operator with
$g(p,k)$ $=e(p)+e(p-k)-2\kappa k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|=\kappa|p|^{2}-\kappa|p-k|^{2}-2\kappa k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|$
$=\kappa\left(2k\cdot(p-k_{\textnormal{F}}\hat{\omega}_{\alpha})-|k|^{2}\right)$
where we used $e(p)$ as defined in (2.8). The bound follows from [BNP+21b, Lemma 5.4] which depends on $\|g\|_{l^{\infty}}$. This can be estimated by using $\text{diam}(B_{\alpha})\leq CN^{\frac{1}{3}}M^{-\frac{1}{2}}$ such that $|g(p,k)|\leq C\kappa N^{\frac{1}{3}}M^{-\frac{1}{2}}$. ∎
@@
It holds for $k\in\Gamma^{\textnormal{nor}}$, $\alpha\in\mathcal{I}_{k}$ and all $\psi\in\mathcal{F}$
$[\mathbb{D}_{\textnormal{B}},c_{\alpha}^{\ast}(k)]=2\kappa k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)^{\ast}$
@@
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|^{2}$ $\leq C\left(\kappa k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\right)^{2}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|^{2},$
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|$ $\leq C\kappa k_{\textnormal{F}}M^{\frac{3}{2}}N^{-\frac{2}{3}+\delta}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|,$
@@
It holds $\mathfrak{E}_{\alpha}^{\text{B}}(k)=2\kappa k_{\textnormal{F}}\sum_{l}|l\cdot\hat{\omega}_{\alpha}|\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\chi(\alpha\in\mathcal{I}_{l})$ and therefore
@@
$\leq\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma^{\textnormal{nor}}}||2\kappa k_{\textnormal{F}}|l\cdot\hat{\omega}_{\alpha}|\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\psi\|\bigg{)}^{2}\leq C\kappa k_{\textnormal{F}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma^{\textnormal{nor}}}||\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\psi\|\bigg{)}^{2}$
$\leq C\kappa k_{\textnormal{F}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma^{\textnormal{nor}}}CMN^{-\frac{2}{3}+\delta}||\mathcal{N}c_{\alpha}(l)\psi\|\bigg{)}^{2}$
$\leq C\kappa k_{\textnormal{F}}\left(CMN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\sum_{l\in\Gamma^{\textnormal{nor}}}\sum_{l^{\prime}\in\Gamma^{\textnormal{nor}}}\|c_{\alpha}(l)(\mathcal{N}-2)\psi\|\ \|c_{\alpha}(l^{\prime})(\mathcal{N}-2)\psi\|$
$\leq C\kappa k_{\textnormal{F}}\left(CMN^{-\frac{2}{3}+\delta}\right)^{2}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|^{2}$
where we used in the second line $l\in\Gamma^{\textnormal{nor}}$ bounded, Lemma A.3 in the third line, $\mathcal{N}c_{\alpha}(l)=c_{\alpha}(l)(\mathcal{N}-2)$ in the fourth line and Cauchy-Schwarz and Lemma A.4 in the last line.
@@
Proof.
We show that the inequality derived in [BNP+21b, eq. (4.10)] holds here independent of our adjustments. The proof in [BNP+21b, eq. (4.10)] makes use of a decomposition of the uncovered Fermi surface
$U\coloneqq\big{(}B_{\textnormal{F}}^{c}\cap(B_{\textnormal{F}}+k)\big{)}\ \backslash\ \bigcup_{\alpha\in\mathcal{I}_{k}}\big{(}B_{\alpha}\cap(B_{\alpha}+k)\big{)}$
into $Y=\{p\in U:e(p)+e(p-k)\leq 4\kappa N^{\frac{1}{3}-\delta}\}$. Although the decomposition depends on the kinetic energy the final bound is independent of the kinetic energy coupling $\kappa$. The condition $e(p)+e(p-k)\leq 4\kappa N^{\frac{1}{3}-\delta}$ is chosen to imply $\hat{p}\cdot\hat{k}\leq CN^{-\delta}$ due to $|p|\sim N^{1/3}$ and $|k|\sim 1$. Thus, one derives $|U\backslash Y|\in\mathcal{O}(N^{1/3}M^{1/2})$ which gives the bound on $U\backslash Y$ with Cauchy-Schwarz. For the bound on $Y$ insert the bound (see [BPSS23, eq. (4.10)])
$\sum_{p\in B_{\textnormal{F}}^{c}\cap(B_{\textnormal{F}}+k)}\frac{1}{e(p)+e(p-k)}\leq C\kappa^{-1}N^{\frac{1}{3}}$
into
$\sum_{p\in B_{\textnormal{F}}^{c}\cap(B_{\textnormal{F}}+k)}1=\sum_{p\in B_{\textnormal{F}}^{c}\cap(B_{\textnormal{F}}+k)}\frac{4\kappa N^{\frac{1}{3}-\delta}}{e(p)+e(p-k)}$
canceling the $\kappa$-dependence. In total we end up with the same inequalities. ∎

### v2

@@ Appendix A Estimates within the bosonization framework
$=2k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)^{\ast}$
One makes the identification $\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\equiv c_{\alpha}^{g}(k)$ representing a weighted operator (see [BNP+21b, eq. (5.11)]) with
$g(p,k)$ $=e(p)+e(p-k)-2k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|=|p|^{2}-|p-k|^{2}-2k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|$
$=\left(2k\cdot(p-k_{\textnormal{F}}\hat{\omega}_{\alpha})-|k|^{2}\right)$
where we used $e(p)$ as defined in (2.8). The bound follows from [BNP+21b, Lemma 5.4] which depends on $\|g\|_{l^{\infty}}$. This can be estimated by using $\text{diam}(B_{\alpha})\leq CN^{\frac{1}{3}}M^{-\frac{1}{2}}$ such that $|g(p,k)|\leq CN^{\frac{1}{3}}M^{-\frac{1}{2}}$. ∎
@@
It holds for $k\in\Gamma$, $\alpha\in\mathcal{I}_{k}$ and all $\psi\in\mathcal{F}$
$[\mathbb{D}_{\textnormal{B}},c_{\alpha}^{\ast}(k)]=2k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{B}}_{\alpha}(k)^{\ast}$
@@
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|^{2}$ $\leq C\left(k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\right)^{2}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|^{2},$
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}_{\alpha}^{\text{B}}(k)\psi\|$ $\leq Ck_{\textnormal{F}}M^{\frac{3}{2}}N^{-\frac{2}{3}+\delta}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|.$
@@
It holds $\mathfrak{E}_{\alpha}^{\text{B}}(k)=2k_{\textnormal{F}}\sum_{l}|l\cdot\hat{\omega}_{\alpha}|\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\chi(\alpha\in\mathcal{I}_{l})$ and therefore
@@
$\leq\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma}||2k_{\textnormal{F}}|l\cdot\hat{\omega}_{\alpha}|\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\psi\|\bigg{)}^{2}\leq Ck_{\textnormal{F}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma}||\mathcal{E}_{\alpha}^{\ast}(l,k)c_{\alpha}(l)\psi\|\bigg{)}^{2}$
$\leq Ck_{\textnormal{F}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma}CMN^{-\frac{2}{3}+\delta}||\mathcal{N}c_{\alpha}(l)\psi\|\bigg{)}^{2}$
$\leq Ck_{\textnormal{F}}\left(CMN^{-\frac{2}{3}+\delta}\right)^{2}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\sum_{l\in\Gamma}\sum_{l^{\prime}\in\Gamma}\|c_{\alpha}(l)(\mathcal{N}-2)\psi\|\ \|c_{\alpha}(l^{\prime})(\mathcal{N}-2)\psi\|$
$\leq Ck_{\textnormal{F}}\left(CMN^{-\frac{2}{3}+\delta}\right)^{2}\|(\mathcal{N}+1)^{\frac{3}{2}}\psi\|^{2}$
where we used in the second line $l\in\Gamma$ bounded, Lemma A.3 in the third line, $\mathcal{N}c_{\alpha}(l)=c_{\alpha}(l)(\mathcal{N}-2)$ in the fourth line and Cauchy-Schwarz and Lemma A.4 in the last line.
@@ Acknowledgment
Declaration
Conflict of interest
The authors declare that there is no conflict of interest.